Woody Woodpecker picked at a 17m wooden pole until it cracked and the upper part fell, reaching a point on the ground 10m from the foot of the pole. Since the pole had not completely broken off, Woody pecked away where it had cracked. How far was Woody above the ground?
Answer by nerdybill(7384)

Let x = distance Woody above ground
then
hypotenuse = 17-x
.
From Pythagorean:
x^2 + 10^2 = (17-x)^2
x^2 + 100 = 289 - 34x + x^2
100 = 289 - 34x
-189 = -34x
-189/-34 = x
5.559m = x